OpenStudy (anonymous):

Solve: x3 – 2x2 – 48x = 0

jhonyy9 (jhonyy9):

x(x^2 -2x -48)=0 x(x+6)(x-8)=0 x_1=0 x+6=0 x_2=-6 x-8=0 x_3=8